OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). a. -9x+27 b. -9x+25 c. -9x+5 d. -9x-2 can someone tell me how to solve this

OpenStudy (sreekar369):

we know g(x)=-x+3 then f(g(X))=f(-x+3) then substitute the x value

OpenStudy (sreekar369):

@chorrigan15

OpenStudy (anonymous):

so it would be b @sreekar369

OpenStudy (sreekar369):

u r \[correct\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait!! i think it's c @chorrigan15 @sreekar369

OpenStudy (anonymous):

\[f(g(x))=(-(9x-2)+3)\]\[=-9x+2+3\]\[=?\]solve it @chorrigan15

OpenStudy (sreekar369):

u r wrong

OpenStudy (anonymous):

why ?

OpenStudy (sreekar369):

g(x)=-x+3 f(g(x))=f(-x+3) how will u substitute it tht reverse

OpenStudy (anonymous):

f(g(x))=-(9x-2)+3

OpenStudy (sreekar369):

=>f(-x+3) tht means u shuld sub the value in f(x) value lol

OpenStudy (anonymous):

Srry Overlook You're rite @sreekar369 f(-x+3) 9(-x+3)-2 -9x+27-2 -9x+5

OpenStudy (anonymous):

-9x+25*

OpenStudy (sreekar369):

ok ok